OpenStudy (anonymous):

Find an equation of the line through (4,-1) AND parallel to x+y=8. Enter the equation in form of y=mx+b

OpenStudy (anonymous):

you would use point slope form y-y1=m(x-x1) where y1, x1 are your given points. m would be 1 because it is parallel. The answer would be y=-x+5

OpenStudy (anonymous):

First you have to put x + y = 8 in y=mx+b form y = 8 -x Does that make sense so far?

OpenStudy (anonymous):

Sorry i meant m=-1, because the line is parallel to x+y=8

OpenStudy (anonymous):

Yes, Thanks.

OpenStudy (anonymous):

You then have to do what Princies said. Use the formula y-y1 = m(x - x1) to find the equation Plug in the point to x and y to solve.

OpenStudy (anonymous):

y + 1 = -1( x - 4)

OpenStudy (anonymous):

The 1 and -1 will be cross canceled. Leading y=-x+5 like @Principes said

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Thanks!